OpenStudy (anonymous):

My next question is if you were going to write a single logarithm (log(5) - log(x) - log(y)) would be log(xy/5)

OpenStudy (anonymous):

\[\log(5)-\log(x)=\log(\frac{5}{x})\] is a start

OpenStudy (anonymous):

you got this?